Subject: RAITH STATUS
From: James Conway <toecutter4ranger@gmail.com>
Date: Fri, 8 Jan 2010 15:06:45 -0800
Fri, 8 Jan 2010 15:06:45 -0800
RAITH STATUS 15:00 hours Friday:

neslab chillers up and running one on line on the system albeit with reduced
flow...
Brought up the IEE-4888 interfaces and they tested fine. Last night's piezo
control errors simply may have been a comunication hang on the IEEE-4888
buss or an channel addressing drop.

The vacuum system quickly pumped down and the FE-Gun stayed up since this
mornings restart at 11. (Miracles do happen!)

Tested loader after belt tension adjustments and it completed one load lock
exchange cycle.  All new robot arm limit settings established to avoid SW
conflicts encountered yesterday.  We are unloading users samples written
last night  and will soon reload more samples and test loader again.  If you
wish to load with us come on down -- we pump at 15:30...

Further loader testing in progress.

We will take the system down again monday abouty 12 - 1 PM to replace the Z
motor in stage and to resolve all other issues with RAITH Filed service.
More engineers are coming to help us through this Monday.

 USERS ARE REQUESTED TO MOVE THEIR CORAL RESERVATIONS TO BEYOND WED. AT
16:00 HOURS.

Thank you for your support!

JWC


RAITH STATUS 15:00 hours Friday:

neslab chillers up and running one on line on the system albeit with reduced flow...
Brought up the IEE-4888 interfaces and they tested fine. Last night's piezo control errors simply may have been a comunication hang on the IEEE-4888 buss or an channel addressing drop.

The vacuum system quickly pumped down and the FE-Gun stayed up since this mornings restart at 11. (Miracles do happen!)

Tested loader after belt tension adjustments and it completed one load lock exchange cycle.  All new robot arm limit settings established to avoid SW conflicts encountered yesterday.  We are unloading users samples written last night  and will soon reload more samples and test loader again.  If you wish to load with us come on down -- we pump at 15:30...

Further loader testing in progress.

We will take the system down again monday abouty 12 - 1 PM to replace the Z motor in stage and to resolve all other issues with RAITH Filed service.  More engineers are coming to help us through this Monday.

 USERS ARE REQUESTED TO MOVE THEIR CORAL RESERVATIONS TO BEYOND WED. AT 16:00 HOURS.

Thank you for your support!

JWC